Problem 18.46

An engine receives 790J of heat from a hot reservoir and gives off 450J of heat to a cold reservoir.

Part A

What is the work done?

ANSWER:

W =   J  

Part B

What is the efficiency of this engine?

ANSWER:

e =
Answer #1

PART-1

W =Qh-Qc

W=790-450

W=340J

------------------------------------------------------------------------------------

PART-2

efficency = (Qh-Qc)/Qh

eff= 790-450/790

efficiency =0.4303

efficiency=43.03%
Answer #2


apply Work Done W = dQ-dU

W = 790-450 =   340 J -<<<<<<<<<<<<,answer

-------------

efficiency e = 1-T2/T1

e = W/Q

e = 340/790

e = 43.03 % effiecint -<<<<<<<<<<<<,answer